# Latent Diffusion Model Based Denoising Receiver for 6G Semantic Communication: From Stochastic Differential Theory to Application

### v1

@@ Abstract
In this paper, a novel semantic communication framework empowered by generative artificial intelligence (GAI) is proposed, specifically leveraging the capabilities of diffusion models (DMs). A rigorous theoretical foundation is established based on stochastic differential equations (SDEs), which elucidates the denoising properties of DMs in mitigating additive white Gaussian noise (AWGN) in latent semantic representations. Crucially, a closed-form analytical relationship between the signal-to-noise ratio (SNR) and the denoising timestep is derived, enabling the optimal selection of diffusion parameters for any given channel condition. To address the distribution mismatch between the received signal and the DM’s training data, a mathematically principled scaling mechanism is introduced, ensuring robust performance across a wide range of SNRs without requiring model fine-tuning. Built upon this theoretical insight, we develop a latent diffusion model (LDM)-based semantic transceiver, wherein a variational autoencoder (VAE) is employed for efficient semantic compression, and a pretrained DM serves as a universal denoiser. Notably, the proposed architecture is fully training-free at inference time, offering high modularity and compatibility with large-scale pretrained LDMs. This design inherently supports zero-shot generalization and mitigates the challenges posed by out-of-distribution inputs. Extensive experimental evaluations demonstrate that the proposed framework significantly outperforms conventional neural-network-based semantic communication baselines, particularly under low SNR conditions and distributional shifts, thereby establishing a promising direction for GAI-driven robust semantic transmission in future 6G systems.
@@ I Introduction
The advent of sixth-generation (6G) wireless communication marks a fundamental paradigm shift from conventional bit-level transmission toward semantic-level information exchange [1]. Different from traditional systems that aim to faithfully reconstruct raw data at the receiver, semantic communication seeks to extract, compress, and transmit the underlying meaning of the data to fulfill the intent of the communication task [2]. This shift is primarily driven by the growing demand for intelligent, efficient, and context-aware communication in emerging application scenarios. For instance, in extended reality (XR), the perceived user experience depends more on the accurate reconstruction of spatial semantics than on the pixel-wise fidelity of data [3]. In industrial Internet of Things (IIoT) environments, vast amounts of sensor data are generated continuously, yet only a small fraction of this data holds decision-critical value [4]. Similarly, digital twin systems require real-time synchronization between physical entities and their virtual counterparts, necessitating communication mechanisms that are capable of semantic abstraction and selective information delivery [5]. However, current wireless communication architectures, predominantly based on Shannon’s information theory, are designed to ensure the reliable delivery of symbol sequences rather than understanding their semantic content. While effective in maximizing data throughput under bandwidth constraints, these systems are inherently incapable of reasoning about the contextual importance or task relevance of transmitted information [6]. As spectrum resources become increasingly scarce and the volume of data continues to grow, the limitations of bit-centric models become more pronounced. Consequently, both academia and industry are actively exploring new frameworks for semantic communication that can overcome these fundamental bottlenecks. By rethinking the role of the transmitter and receiver as semantic encoders and decoders, future communication systems must go beyond the Shannon limit and prioritize the transmission of meaningful content over syntactic accuracy [7].
@@
Recent advances in deep learning have led to the proliferation of neural network-based semantic communication frameworks, where the end-to-end system is typically implemented as an autoencoder architecture comprising a neural encoder, a noisy channel, and a neural decoder [8]. These systems demonstrate impressive performance in compressing and reconstructing semantic features, enabling tasks such as image transmission with drastically reduced bit rates. However, despite the promising capabilities of neural network-based semantic communication systems, they exhibit significant limitations in terms of robustness and generalization. One of the most critical challenges arises from their acute sensitivity to signal-to-noise ratio (SNR) variation [9]. Specifically, the semantic features produced by the encoder are susceptible to corruption by additive noise introduced by the wireless channel, resulting in received representations at the decoder that diverge markedly from those observed during training. This leads to pronounced performance degradation under low-SNR or unexpected channel conditions [10]. Fundamentally, this vulnerability stems from a distribution mismatch between training and inference inputs, a phenomenon commonly referred to as the out-of-distribution (OOD) problem [11]. Moreover, the issue is further exacerbated by the intrinsic properties of latent semantic representations. Unlike explicit modalities such as images or audio, latent features are highly abstract, compactly encoded, and often reside in non-Euclidean manifolds [11]. This compressed structure renders them particularly fragile to perturbations. Even minor Gaussian noise interference can induce significant distributional shifts that compromise the underlying semantic structure. Consequently, neural decoders struggle to accurately reconstruct the original content, as they fail to extract meaningful patterns from perturbed latent vectors. This fragility of the latent space poses a unique challenge to semantic communication systems, namely, how to effectively model, diagnose, and mitigate noise-induced performance degradation in compressed semantic domains to improve the fidelity and reliability of the semantic feature decoding. Furthermore, current neural semantic systems are often tailored to data from a particular distribution, limiting their scalability to broader domains. When the distribution of transmitted data changes, even slightly, from the distribution of cats to the distribution of dogs, the performance of the system can deteriorate substantially, necessitating full retraining [2]. Such dependence on narrowly scoped training regimes not only impairs scalability but also introduces a prohibitive computational burden when deployed in dynamic or heterogeneous environments. Consequently, the lack of noise resilience, data generalization, and model adaptability remains a fundamental bottleneck that hinders the widespread application of neural network-based semantic communication in real-world 6G scenarios [12].
The aforementioned limitations arise from a deeper theoretical mismatch between the intrinsic nature of semantic communication and the architectural foundations of current neural network-based solutions [11]. Traditional semantic communication frameworks predominantly adopt autoencoder-based architectures, where the neural backbone is trained in a discriminative manner. While effective in encoding data representations under matched training and inference conditions, such models exhibit a critical limitation: they lack the capacity to generate meaningful outputs from null or severely degraded inputs. This behavior is indicative of a discriminative modeling approach, which aims to extract discriminative features rather than to model the full data distribution. According to statistical learning theory, discriminative models are inherently restricted to learning decision boundaries or conditional mappings, rather than capturing the underlying generative process of the data. However, from an information-theoretic perspective, semantic communication aims to exploit the prior distribution of source data, which is the semantic information, to optimize source and channel encoding jointly. This joint design enables more efficient data compression and enhanced resilience against channel impairments. Crucially, without effective utilization of the data’s prior distribution, it is fundamentally difficult to realize the theoretical performance gains of semantic communication. This insight motivates a paradigm shift from discriminative to generative modeling as the backbone architecture of semantic communication systems. Generative models, by construction, capture the underlying data distribution and therefore enable maximum a posteriori (MAP) estimation of corrupted semantic features, facilitating robust recovery of the original content from noisy observations. Although recent studies have proposed generative-model-based semantic communication frameworks, they often rely on conditional generative models that reconstruct data from transmitted semantic features. However, this approach suffers from two notable drawbacks. First, conditional generative models typically involve manually injected Gaussian noise during training to facilitate sampling. This artificial injection often leads to a mismatch between generated samples and the true distribution of the original data, especially under noisy channel conditions. Second, these models generally depend on conditional encoders to extract semantic features. As previously noted, such encoders are essentially discriminative in nature, and are therefore vulnerable to OOD degradation when exposed to distorted features resulting from channel noise. To address these limitations, a new generative-model-driven framework is required—one that both eliminates the reliance on fragile conditional structures and fully leverages the inherent data priors learned by the generative model. Such a design would enable effective denoising and semantic reconstruction even under adverse channel conditions, thereby enhancing both the robustness and theoretical fidelity of semantic communication systems.
To address the above challenges, this work introduces a novel semantic communication architecture based on latent diffusion models (LDMs) [13]. The key innovation lies in leveraging the generative capability of LDMs to perform semantic-level denoising in the latent space of a variational autoencoder (VAE) framework [14]. Instead of directly operating in the high-dimensional signal domain, our method encodes semantic information into a compact latent representation, which is then transmitted over a noisy channel. Upon reception, the system applies an LDM to iteratively remove additive Gaussian noise from the corrupted latent features through a learned reverse diffusion process. This design enables robust semantic recovery even under severe SNR degradation, without requiring retraining for each noise condition. The theoretical foundation of this architecture is grounded in stochastic differential equations (SDE) theory, which formally describes the diffusion and denoising process as a continuous-time Markov chain driven by Brownian motion [15, 16]. In this framework, we derive a closed-form relationship between the SNR of the received feature map and the appropriate denoising timestep in the LDM. This mapping allows the receiver to adaptively select the number of reverse diffusion steps based on the observed channel conditions, ensuring that the generative denoising process is optimally matched to the level of corruption in the signal. Furthermore, to mitigate the OOD problem caused by mismatched input statistics, we introduce a linear scaling mechanism that transforms the received noisy latent features to align with the prior distribution assumed during the diffusion model’s training phase. The combination of adaptive timestep selection and input scaling endows the proposed architecture with strong generalization capabilities across a wide range of channel noise conditions. Crucially, since the LDM is decoupled from the encoder-decoder training pipeline and used solely for noise suppression, pre-trained diffusion models from general computer vision tasks can be directly integrated without fine-tuning. This plug-and-play property significantly reduces system design complexity and enables continual evolution of the communication system as more powerful generative models become available. By bridging stochastic process theory with semantic communication design, our approach provides both theoretical guarantees and practical robustness, paving the way for scalable, noise-resilient 6G communication systems. The main contributions of this paper are summarized as follows.
  1. 1.
    A novel theoretical foundation for semantic communication is established in this paper by leveraging SDEs to model the relationship between channel-induced Gaussian noise and the reverse diffusion process in generative models. We derive a closed-form analytical expression linking the signal-to-noise ratio (SNR) of the received semantic representation to the appropriate denoising timestep in the diffusion model. This formulation enables precise and adaptive noise mitigation without retraining, offering a mathematically grounded mechanism to handle arbitrary noise levels. The proposed theory not only bridges the gap between stochastic process modeling and neural semantic communication but also enhances robustness under out-of-distribution channel conditions, a longstanding challenge in the field.
  2. 2.
    Building on the above theoretical framework, we propose an end-to-end semantic communication system that integrates a VAE with an LDM. In this architecture, the VAE compresses high-dimensional source data into compact semantic latent variables, which are transmitted over a noisy channel and then denoised using a pretrained LDM. By introducing a scaling mechanism to align the noisy latent features with the training distribution of the LDM, the system effectively mitigates OOD issues. Crucially, the LDM operates independently of the encoder-decoder pipeline and can be replaced by any pretrained diffusion model, enabling zero-shot generalization across diverse SNRs and data types without task-specific fine-tuning.
  3. 3.
    Experimenets
II Related Works and Preliminaries
@@ II-A Semantic Communications
Semantic communication has emerged as a transformative paradigm in the evolution of next-generation wireless systems, particularly 6G, aiming to bridge the gap between signal transmission and meaning understanding. Unlike conventional Shannon-centric systems that prioritize bit-level fidelity, semantic communication seeks to transmit the intended meaning of messages effectively, thus aligning the communication process with the task-specific utility at the receiver. This paradigm shift has triggered intensive theoretical and practical exploration in recent years, resulting in a rich body of literature.
@@
Foundational works such as those by [6, 17] established the conceptual framework for semantic communication, delineating the limitations of Shannon theory in capturing semantic and effectiveness layers of communication, and advocating for AI-integrated approaches. Subsequent surveys, including those by [18] and [19], have categorized enabling technologies such as knowledge-based reasoning, deep learning-based encoders, and task-oriented transmission frameworks. These reviews have also formalized key metrics including semantic entropy, semantic rate-distortion, and semantic channel capacity, marking a theoretical extension beyond classical information theory.
@@
In parallel, significant efforts have been made toward system-level innovations. [20] and [21] reviewed emerging architectures that integrate semantic communication with deep joint source-channel coding (JSCC), generative AI , and federated learning. These contributions underline the role of large language models (LLMs) and semantic knowledge bases (KBs) in enhancing interpretability, compressibility, and adaptability of semantic encoders across modalities such as text, image, audio, and video. The work by Lema et al. [22] highlights the convergence of semantic communication with edge computing and proposes scalable architectures for real-time, context-aware applications.
@@
A number of domain-specific studies have investigated semantic communication across critical verticals. For instance, [23] examines its integration into vehicular networks (V2X), focusing on multimodal semantic extraction, cooperative coding, and dynamic resource allocation. Similarly, other works explore semantic transmission in metaverse platforms, Internet of Things (IoT), wireless sensor networks (WSN), and healthcare, emphasizing the ability to reduce bandwidth consumption and improve robustness through semantic compression and prediction. Despite these advances, several open challenges remain. Theoretical gaps persist in defining a unified framework for semantic information quantification and in establishing universal semantic performance metrics. Additionally, the issues of dynamic KB synchronization, adversarial robustness, interpretability of deep semantic models, and semantic-level security threats have become critical research frontiers. Recent studies, such as those by [24], have introduced performance modeling techniques, such as the Alpha-Beta-Gamma (ABG) expression, to bridge empirical deep learning performance with classical SNR-based communication analysis.
@@ II-B Diffusion Model
Diffusion models are a class of generative models that produce data samples by iteratively reversing a predefined noise-injection process. Originally proposed as an alternative to generative adversarial networks (GANs) [25], diffusion models have demonstrated remarkable performance in diverse generative tasks, including image synthesis, natural language modeling, and structured data generation [26, 27]. Beyond generation, they have also shown promise in perception-oriented applications such as image segmentation, object detection, and model-based reinforcement learning [13]. At their core, diffusion models define a two-stage process: a forward diffusion stage that incrementally perturbs input data with Gaussian noise, and a reverse denoising stage that reconstructs the data through a learned Markov process. Let $\bm{x}_{0}$ denote the clean input. The forward process generates latent variables $\bm{x}_{1},\ldots,\bm{x}_{T}$ via a Markov chain with Gaussian transitions as follows.
$q\left(\bm{x}_{t}\mid\bm{x}_{t-1}\right)=\mathcal{N}\left(\sqrt{1-\beta_{t}},\bm{x}_{t-1},\beta_{t}\bm{I}\right),$ (1)
where $\beta_{t}\in(0,1)$ controls the noise variance at step $t$. By defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, one obtains the closed-form marginal as follows.
$q(\bm{x}_{t}\mid\bm{x}_{0})=\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}},\bm{x}_{0},(1-\bar{\alpha}_{t})\bm{I}\right),$ (2)
Applying the sampling rule, the following align can be obtained.
$\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}},\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}},\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3)
@@
where $\bm{\mu}_{\bm{\theta}}$ is a neural network trained to predict the denoised signal at each step. Sampling proceeds from $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ down to $\bm{x}0$, with each denoising step computed as follows.
$\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}t}}\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)\right)+\beta_{t}\bm{z},$ (5)
The final noise term ensures that the variance of the denoised sample aligns with that of the forward process. While a more precise formulation would scale this term using $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$, empirical studies [28] have shown that using $\beta_{t}$ directly provides a favorable trade-off between computational efficiency and denoising performance.

### v3

@@ Abstract
In this paper, a novel semantic communication framework empowered by generative artificial intelligence (GAI) is proposed, to enhance robustness against both channel noise and transmission data distribution shifts. A theoretical foundation is established using stochastic differential equations (SDEs), from which a closed-form mapping between any signal-to-noise ratio (SNR) and the optimal denoising timestep is derived. Moreover, to address distribution mismatch, a mathematical scaling method is introduced to align received semantic features with the GAI training distribution. Built on this theoretical foundation, a latent diffusion model (LDM)-based semantic communication framework is proposed to combine a variational autoencoder for semantic features extraction, where a pretrained diffusion model is used for denoising. The proposed system is a training-free framework that supports zero-shot generalization, and achieves superior performance under low-SNR and out-of-distribution conditions, offering a scalable and robust solution for future 6G semantic communication systems. Experimental results demonstrate that the proposed semantic communication framework achieves state-of-the-art performance in both pixel-level accuracy and semantic perceptual quality, consistently outperforming baselines across a wide range of SNRs and data distributions without any fine-tuning or post-training.
@@ I Introduction
The research of sixth-generation (6G) wireless communication has shown a growing paradigm shift from conventional bit-level transmission toward semantic-level information exchange [1]. Unlike traditional systems that aim to reconstruct raw data, semantic communication focuses on extracting and transmitting the intended meaning of the data to support task execution [2]. This evolution is driven by increasing demands for intelligent, efficient, and context-aware communication in emerging applications. For example, in extended reality (XR), user experience depends more on spatial semantics than pixel fidelity [3]; in industrial IoT (IIoT), only a small portion of sensor data is decision-critical [4]; and digital twin systems require semantic-level synchronization between physical and virtual entities [5]. However, current systems based on Shannon’s theory focus on reliable symbol delivery and lack the capability to assess contextual or task relevance [6]. Therefore, both academia and industry communities are actively exploring semantic communication frameworks that transcend the Shannon paradigm by prioritizing meaningful content over syntactic precision [7].
Recent advances in deep learning have led to the proliferation of neural network-based semantic communication frameworks, where the end-to-end system is typically implemented as an autoencoder architecture comprising a neural encoder, a noisy channel, and a neural decoder [8]. These systems demonstrate impressive performance in compressing and reconstructing semantic features, enabling tasks such as image transmission with drastically reduced bit rates [9]. However, despite the promising capabilities of neural network-based semantic communication systems, they exhibit significant limitations in terms of robustness and generalization. One of the most critical challenges arises from their acute sensitivity to signal-to-noise ratio (SNR) variations [10]. Specifically, the semantic features produced by the encoder are susceptible to corruption by additive noise introduced by the wireless channel, resulting in the received representations at the decoder diverging markedly from those observed during training. This leads to significant performance degradation under low-SNR or unexpected channel conditions [11]. Fundamentally, this challenge arises from a distribution mismatch between training and inference inputs, a phenomenon commonly referred to as the out-of-distribution (OOD) problem [12]. The issue is further exacerbated by the intrinsic properties of latent semantic representations. Unlike explicit modalities such as images or audio, latent features are highly abstract, compactly encoded, and often reside in non-Euclidean manifolds [12]. This compressed structure renders them particularly fragile to perturbations. Even slight Gaussian noise can induce significant distributional shifts that compromise the underlying semantic structure. Consequently, neural decoders struggle to accurately reconstruct the original content, as they fail to extract meaningful patterns from perturbed latent vectors [11]. This fragility of the latent space poses a huge challenge to semantic communication systems, namely, how to effectively model, diagnose, and mitigate noise-induced performance degradation in compressed semantic domains to improve the fidelity and reliability of the semantic feature decoding.
Furthermore, current neural semantic systems are often trained on data from a particular distribution, limiting their scalability to broader domains. When the distribution of transmitted data changes, even slightly, from the distribution of cats to the distribution of dogs, the performance of the system may deteriorate substantially, necessitating full retraining [2]. Such dependence on narrowly scoped training regimes not only impairs scalability but also introduces a prohibitive computational burden when deployed in dynamic or heterogeneous environments. Consequently, the lack of noise resilience, data generalization, and model adaptability remains a fundamental bottleneck that hinders the widespread application of neural network-based semantic communication in practical 6G scenarios [13].
The above limitations arise from a fundamental mismatch between the nature of semantic communication and the predominantly discriminative architectures used in current neural solutions [12]. Most systems adopt auto-encoder backbones, training neural networks to extract semantic features under matched training and inference conditions. Although these models perform adequately on in-distribution data, they deteriorate sharply when the channel is highly noisy or when test statistics deviate from the training distribution [14]. From an information-theoretic perspective, the joint source–channel coding should leverage the data’s prior distribution to achieve superior compression and robustness [15]. However, discriminative networks are ill-equipped to learn full data priors. This motivates a shift toward generative modelling, which, in principle, captures the underlying distribution and enables maximum a posteriori recovery from corrupted latent features. Recent attempts to incorporate generative models have relied on conditional diffusion or GAN architectures [16, 9], yet these approaches inject artificial Gaussian noise during training, which creates a mismatch with real channels, and still depend on discriminative conditional encoders, leaving them vulnerable to OOD inputs. Consequently, a fully generative framework that eliminates conditional bottlenecks and directly exploits learned priors is essential for robust denoising and semantic reconstruction across diverse channel conditions.
To address the aforementioned challenges, in this paper, we propose a novel semantic communication architecture based on latent diffusion models (LDMs) [17]. The core idea is to harness the generative power of LDMs for semantic-level denoising in the latent space of a variational autoencoder (VAE) [18]. Rather than operating in the high-dimensional signal domain, the proposed method encodes semantic information into a compact latent vector, transmits it over a noisy channel, and applies a pretrained LDM at the receiver to iteratively remove Gaussian noise via reverse diffusion. This enables robust semantic recovery under severe SNR degradation without retraining. The framework is theoretically grounded in stochastic differential equation (SDE) theory, modeling diffusion as a continuous-time Markov process [19, 20]. A closed-form mapping between the SNR and the diffusion timestep is derived, allowing the receiver to adaptively select the number of denoising steps based on channel conditions. To mitigate the OOD issue from mismatched input statistics, a linear scaling mechanism is introduced to align received features with the training distribution of the LDM. This dual adaptation strategy ensures strong generalization across diverse noise levels. Since the LDM operates independently of the encoder–decoder pipeline and serves solely as a denoiser, pretrained models from vision tasks can be seamlessly integrated without fine-tuning. This plug-and-play capability reduces system complexity and supports continual performance enhancement as more advanced generative models emerge. The main contributions of this paper are summarized as follows.
  1. 1.
    Based on SDE, a novel semantic communication theory is proposed, which models the relationship between channel-induced Gaussian noise and the reverse diffusion process using stochastic differential equations. A closed-form mapping between the SNR and the optimal denoising timestep is derived, enabling adaptive noise suppression without retraining, and addressing the long-standing challenge of OOD robustness in semantic communication.
  2. 2.
    Built on the proposed semantic communication theory, an LDM-based framework is presented, which integrates a VAE for semantic compression, a scaling mechanism for distribution alignment, and a pretrained LDM for denoising in latent space. As the LDM operates independently of the encoder–decoder pair, it supports seamless substitution with a wide range of pretrained diffusion models, significantly simplifying system deployment and upgrade.
  3. 3.
    Experiment results show the proposed semantic communication framework achieves SOTA performance in both pixel-level accuracy and semantic perceptual quality, across a wide range of SNR and transmission data distribution without any fine-tuning or post-training.
@@
II Preliminaries and Related Works
@@ II-A Semantic Communications
Semantic communication has emerged as a transformative paradigm in the evolution of next-generation wireless systems, aiming to bridge the gap between signal transmission and meaning understanding. Unlike conventional Shannon-centric systems that prioritize bit-level fidelity, semantic communication seeks to effectively transmit the intended meaning of messages, thus aligning the communication process with the task-specific utility at the receiver. This paradigm shift has triggered intensive theoretical and practical exploration in recent years, resulting in a rich body of literature.
@@
Foundational works such as [6] and [21] established the conceptual framework for semantic communication, delineating the limitations of Shannon theory in capturing semantic and effectiveness layers of communication, and advocating for AI-integrated approaches. Subsequent surveys, including [22] and [9], have categorized enabling technologies such as knowledge-based reasoning, deep learning-based encoders, and task-oriented transmission frameworks. These reviews have also formalized key metrics, including semantic entropy, semantic rate-distortion, and semantic channel capacity, marking a theoretical extension beyond classical information theory.
@@
In parallel, significant efforts have been made toward system-level innovations. In [23] and [24], the authors review emerging architectures that integrate semantic communication with deep joint source-channel coding (JSCC), generative AI, and federated learning. These contributions underline the role of large language models (LLMs) and semantic knowledge bases (KBs) in enhancing interpretability, compressibility, and adaptability of semantic encoders across modalities such as text, image, audio, and video. The work by Lema et al. [25] highlights the convergence of semantic communication with edge computing and proposes scalable architectures for real-time, context-aware applications.
@@
A number of domain-specific studies have investigated semantic communication across critical verticals. For instance, Ye et al. examine its integration into vehicular networks, focusing on multimodal semantic extraction, cooperative coding, and dynamic resource allocation [26]. Similarly, other works explore semantic transmission in metaverse platforms, Internet of Things (IoT), wireless sensor networks (WSN), and healthcare, emphasizing the ability to reduce bandwidth consumption and improve robustness through semantic compression and prediction. Despite these advances, several open challenges remain. Theoretical gaps persist in defining a unified framework for semantic information quantification and in establishing universal semantic performance metrics. Additionally, the issues of dynamic KB synchronization, adversarial robustness, interpretability of deep semantic models, and semantic-level security threats have become critical research frontiers. Recent studies, includes [27], have introduced performance modeling techniques, such as the Alpha-Beta-Gamma (ABG) expression, to bridge empirical deep learning performance with classical SNR-based communication analysis.
@@ II-B Diffusion Model
Diffusion models are a class of generative models that produce data samples by iteratively reversing a predefined noise-injection process. Originally proposed as an alternative to generative adversarial networks (GANs) [28], diffusion models have demonstrated remarkable performance in diverse generative tasks, including image synthesis, natural language modeling, and structured data generation [29, 30]. Beyond generation, they have also shown promise in perception-oriented applications such as image segmentation, object detection, and model-based reinforcement learning [17, 31, 32]. At their core, diffusion models define a two-stage process: a forward diffusion stage that incrementally perturbs input data with Gaussian noise, and a reverse denoising stage that reconstructs the data through a learned Markov process. Let $\bm{x}_{0}$ denote the clean input. The forward process generates latent variables $\bm{x}_{1},\ldots,\bm{x}_{T}$ via a Markov chain with Gaussian transitions given by
$q\left(\bm{x}_{t}\mid\bm{x}_{t-1}\right)=\mathcal{N}\left(\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I}\right),$ (1)
where $\mathcal{N}(\bm{\mu},\sigma^{2}\bm{I})$, is the Gaussian distribution with the mean if $\bm{\mu}$ and covariance of $\sigma\bm{I}$, $\bm{I}$ is an identical matrix, and $\beta_{t}\in(0,1)$ controls the noise variance at step $t$. By defining $a_{t}=1-\beta_{t}$ and $\bar{a}_{t}=\prod_{s=1}^{t}a_{s}$, one obtains the closed-form marginal as follows.
$q(\bm{x}_{t}\mid\bm{x}_{0})=\mathcal{N}\left(\sqrt{\bar{a}_{t}}\bm{x}_{0},(1-\bar{a}_{t})\bm{I}\right),$ (2)
Applying the sampling rule, the following equation can be obtained.
$\bm{x}_{t}=\sqrt{\bar{a}_{t}}\bm{x}_{0}+\sqrt{1-\bar{a}_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}).$ (3)
@@
where $\bm{\mu}_{\bm{\theta}}$ is a neural network trained to predict the denoised signal at each step. Sampling proceeds from $\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ down to $\bm{x}_{0}$, with each denoising step computed as follows.
$\bm{x}_{t-1}=\frac{1}{\sqrt{a_{t}}}\left(\bm{x}_{t}-\frac{1-a_{t}}{\sqrt{1-\bar{\alpha}t}}\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t)\right)+\beta_{t}\tilde{\bm{\epsilon}},$ (5)
$\tilde{\bm{\epsilon}}\sim\mathcal{N}(\bm{0},\bm{I})$ (6)
where $\tilde{\bm{\epsilon}}$ is the Gaussian noise added in the reverse process. The final noise term ensures that the variance of the denoised sample aligns with that of the forward process. While a more precise formulation would scale this term using $\tilde{\beta}_{t}=\frac{1-\bar{a}_{t-1}}{1-\bar{a}_{t}}\beta_{t}$, empirical studies [33] have shown that using $\beta_{t}$ directly provides a favorable trade-off between computational efficiency and denoising performance.